b. P(-1.69 SZ SZ) = 0.9545 C. P(Z > 20) = 0.0048 d. P(Z, SZS 2.00) = 0.1359 e. P(Z > Z) = 0.9713 f. PC-Z, SZSZ) = 0.2358
Answer #1

Solution:

b)

P( -1.69 \leq Z \leq z0 ) = 0.9545

P( Z \leq z0 ) = 0.9545 + P( Z \leq -1.69 )

P( Z \leq z0 ) = 0.9545 + 0.025

P( Z \leq 2.044 ) =0.9795

z0 = 2.044

phpfhTUiC.png

c)

P( Z < z0 ) = 0.9952

P( Z < 2.59 ) = 0.9952

z0 = 2.59

phpyb8rmu.png

d)

P( Z < z0 ) = 0.8413

P( Z < 1) = 0.8413

z0 = 1

phpz2RdBM.png

e)

P( Z < z0 ) = 0.0287

z0 = -1.9

phpCPs5t1.png

f)

P( Z < z0 ) = 0.6179

z0 = -0.87

and

z0 = 0.87

phppd5Dc5.png